A businessman makes the St. Louis to Chicago (Round Trip) trip 5 times a year in a car which gets 24 mpg. What would the fuel cost be for these 5 journeys if gas costs $2.50 per gallon for all 5 journeys?
Answer by nerdybill(7384)

A businessman makes the St. Louis to Chicago (Round Trip) trip 5 times a year in a car which gets 24 mpg. What would the fuel cost be for these 5 journeys if gas costs $2.50 per gallon for all 5 journeys?
.
First, you need to determine the distance between the two locations.
Distance is 415 kilometers or 258 miles
Ref:http://www.timeanddate.com/worldclock/distanceresult.html?p1=605&p2=64
.
Round trip:
2*258 = 516 miles
.
Distance for 5 round trips:
5*516 = 2580 miles
.
Gallons used:
2580/24 = 107.5
.
Fuel costs:
2.50*107.5 = $268.75